Suppose that shopping times for customers at a local mall are normally distributed with a
known population standard deviation of 20 minutes. A random sample of 64 shoppers in the local
grocery store had a mean time of 75 minutes. Find the standard error, margin of error, and the
upper and lower confidence limit of a 95% confidence interval for the population mean (
1
Expert's answer
2021-07-16T03:26:32-0400

"\sigma=20 \\

n=64 \\

\bar{x}=75"

The value of z at 95% confidence level is 1.96.

The standard error is

"SE = \frac{\sigma}{\sqrt{n}} \\

= \frac{20}{\sqrt{64}}=2.5"

The margin of error is

"ME = Z \times SE \\

= 1.96 \times 2.5 = 4.9"

And the 95% confidence interval is

"\bar{x}±ME = 75 ± 4.9 \\

=(70.1, 79,9)"